# Cohomology of the pure symmetric automorphisms of right-angled Artin groups

### 15 Apr 2026 v1

@@ Abstract.
We compute the cohomology groups of the pure symmetric outer automorphism group $\Sigma\mathrm{POut}(A_{\Gamma})$ and the pure symmetric automorphism group $\Sigma\mathrm{PAut}(A_{\Gamma})$ of a right-angled Artin group $A_{\Gamma}$. Using the equivariant spectral sequence arising from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on the generalized McCullough-Miller complex $\mathrm{MM}_{\Gamma}$, we show that $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is free abelian and we compute its rank in terms of the combinatorics of certain poset. Applying the Lyndon-Hochschild-Serre spectral sequence and the Leray-Hirsch theorem we do the same for $H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))$. In both cases the cohomology algebra is generated in degree 1. Finally, we introduce the Generalized Brownstein-Lee Conjecture, proposing a presentation of $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma})$, and prove that it holds in dimension $2$.
@@ 1. Introduction
In [15], McCullough and Miller constructed a contractible simplicial complex $\mathrm{MM}_{n}$ that admits an action of the group $\Sigma\mathrm{POut}(F_{n})$ with free abelian stabilizers. The construction is based on the Whitehead poset $\mathrm{Wh}_{n}$, whose elements can be expressed as labeled trees. In [9], Jensen, McCammond and Meier use the McCullough-Miller complex to compute $H^{*}\left(\Sigma\mathrm{POut}(F_{n})\right)$ and $H^{*}\left(\Sigma\mathrm{PAut}(F_{n})\right)$, the cohomology groups of the (outer) pure symmetric automorphisms of the free group; evenmore, they prove the Brownstein-Lee conjecture [5], which stablish a presentation of the cohomology ring $H^{*}\left(\Sigma\mathrm{PAut}(F_{n})\right)$.
In [1], both the Whiteahead poset and the construction of McCullough and Miller are generalized from the case of $\Sigma\mathrm{POut}(F_{n})$ to $\Sigma\mathrm{POut}({A_{\Gamma}})$. This new complex, $\mathrm{MM}_{\Gamma}$, admits an action of the group $\Sigma\mathrm{POut}({A_{\Gamma}})$ with free abelian stabilizers and it is also contractible. The powerful combinatorial machinery derived from the labeled trees in $\mathrm{Wh}_{n}$ is the key to prove the main results in [9]. However, vertices in $\mathrm{Wh}_{\Gamma}$ cannot be expressed as labeled trees in general, only as families of based partitions, so most arguments need to be adapted.
In this paper we will generalize some combinatorial results from $\mathrm{Wh}_{n}$ to $\mathrm{Wh}_{\Gamma}$ and use the equivariant spectral sequence derived from the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action over $\mathrm{MM}_{\Gamma}$ to compute the cohomology groups $H^{q}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ and $H^{q}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$. We show (see Sections 2 and 4 for definitions):
@@
Besides, the cohomology ring $H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$, where the $C_{I}^{j}$ are canonical automorphisms.
@@
$H^{q}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}.$
Besides, the cohomology ring $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$ is generated in degree 1.
Furthermore, we generalize the relations from the Browstein-Lee Conjecture to the case of RAAGs, formulate the Generalized Browstein-Lee Conjecture, and prove that it holds in dimension 2.
Conjecture A (Generalized Browstein-Lee Conjecture).
Let $\gamma_{A}^{i}$ be a one dimensional class in $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$, where $A$ is a connected component in $\Gamma-\mathrm{st}(v_{i})$. For each pair $v_{i},v_{j}$ of two non-adjacent vertices in $\Gamma$, let $D^{j},D^{i}$ be the respective dominant components and $\mathcal{C}$ a shared component. Then $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$ is generated by the one-dimensional classes $\gamma_{A}^{i}$, modulo the following relations for each shared component $\mathcal{C}$ of $v_{i},v_{j}$.
@@
$\phi:H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)\twoheadrightarrow R$
@@
The paper is organized as follows: In Section 2 we explain some basic results about MMΓ, obtained from [1]. In Section 3, we go over the main definitions in [9]. In Section 4 we generalize the notion of one-two automorphism which plays a key role in [9] and do some technical work that will be needed later. In Section 5 we apply the previous results to the spectral sequence derived from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on MMΓ to compute $H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ and prove Theorem A. In Section 6, we use the Lyndon-Hoschild-Serre spectral sequence associated to $1\to A_{\Gamma}\to\Sigma\mathrm{PAut}(A_{\Gamma})\to\Sigma\mathrm{POut}(A_{\Gamma})\to 1$ and the Leray-Hirsch Theorem to compute the $H^{*}\left(\Sigma\mathrm{PAut}(A_{\Gamma})\right)$ groups and prove Theorem B. In Section 7 we take the first steps into proving the Brownstein-Lee conjecture for $\Sigma\mathrm{PAut}(A_{\Gamma})$ and we prove Theorem C.
@@ 2. The McCullough-Miller complex
We will use $\Gamma$ to denote a finite simplicial graph with vertex set $V_{\Gamma}$ and set of edges $E(\Gamma)$. Very often we will identify $\Gamma$ or a subgraph of $\Gamma$ with its vertex set. Let $v\in V_{\Gamma}$, $\mathrm{lk}(v)$ denotes the link of $v$, which is the subgraph induced by the vertices adjacent to $v$. We will define the star of $v$, $\mathrm{st}(v)$, as the subgraph induced by $\mathrm{lk}(v)\cup\{v\}$.
@@ 2.1. Pure symmetric automorphisms of RAAGs
Given $v\in L=V_{\Gamma}\cup V_{\Gamma}^{-1}$ and $A$ a union of connected components of $\Gamma-\mathrm{st}(v)$, the partial conjugation of $A$ by $v$ is the automorphism $C_{A}^{v}$ given by
@@
The group $\Sigma\mathrm{PAut}(A_{\Gamma})$ is the group generated by all partial conjugations $C_{A}^{u}$ for $u\in V_{\Gamma}$ and $A$ a connected component of $V_{\Gamma}-\mathrm{st}(u)$, subject to the following relations:
@@
The nuclear vertex type, denoted $\mathcal{O}$, is the vertex type such that for each $v\in\Gamma$ the associated based partition is trivial, i.e., $\mathcal{O}_{v}=\{\{v\},P\}$, where $P=\Gamma-\mathrm{st}(v)$. It id the unique minimal element in the $\mathrm{Wh}_{\Gamma}$ poset.
@@
We say that a generating system $X$ of $A_{\Gamma}$ is a basis if there is some $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$ such that $X=\alpha(V_{\Gamma})$. A marked based partition is a pair $(X,\tau_{u})$ where $X$ is an ordered basis of $A_{\Gamma}$ and $\tau_{u}$ is a based partition and a marked vertex type $(X,\tau)$ is the family of marked based partitions $(X,\tau_{u})$ where $\tau_{u}$ runs over the set of based partitions of a vertex type $\tau$. By marking a based partition $\tau_{u}$ we understand that we are relabeling the elements in the partition using $X$ (via the automorphism $\alpha$). We extend the partial order of Definition 2.11 to marked vertex types as follows: $(X_{1},\tau)\leq(X_{2},\tau^{\prime})$ if $X_{1}=X_{2}$ and $\tau\leq\tau^{\prime}$. The marked vertex types form a poset and the group $\Sigma\mathrm{PAut}(A_{\Gamma})$ acts on this poset via
@@
where $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$. We say that $\alpha$ is carried by a marked vertex type $(X,\tau)$ if $\gamma\alpha\gamma^{-1}$ is carried by $\tau$ where $\gamma(X)=V_{\Gamma}$.
@@
The definition of the action above implies that the stabilizer $H=\mathrm{Stab}_{\Sigma\mathrm{POut}({A_{\Gamma}})}[V_{\Gamma},\tau]$ of the vertex $V_{\Gamma},\tau]$ under the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action is precisely the subgroup generated by those automorphisms carried by the vertex type $\tau$. This implies that the stabilizer of the vertex $X,\tau]$ is $\alpha H\alpha^{-1}$ where $\alpha$ is such that $X=\alpha(V_{\Gamma})$. It is proved in [1, Theorem A] that these stabilizers are free abelian, that their rank can be computed in terms of the length of the based partitions of $\tau$ and that they yield in fact all the cell stabilizers of the McCullough-Miller complex $\mathrm{MM}_{\Gamma}$.
@@
$\{\mathrm{Stab}_{\Sigma\mathrm{POut}(A_{\Gamma})}[V_{\Gamma},\tau]\mid\tau\text{ vertex type}\}$
@@ 4. Choosing the basis for $\Sigma\mathrm{POut}({A_{\Gamma}})$
From now on we fix an order in the elements in $V_{\Gamma}$, and denote them by $v_{1},...,v_{n}$. Sometimes, for simplicity and in the aim of a clear notation, we will just identify $v_{i}$ with $i$; in this sense, we will often put $\tau_{i}$ to denote the based partition $\tau_{v_{i}}$, or $C^{j}_{I}$ for the partial conjugation $C^{v_{j}}_{I}$. Let $\tau$ be a vertex type in MMΓ and $\tau_{i}$ a based partition of $\tau$. The product of the partial conjugations associated to each petal of $\tau_{i}$ gives the inner automorphism associated to $v_{i}$, so it is trivial in the outer group. We need to have a way to choose families of partial conjugations associated to $\tau_{i}$ which are independent.
@@
Following the definition, we have that if $\bar{C}_{I_{1}}^{i},...,\bar{C}_{I_{k}}^{i}$ is the set of canonical outer-automorphisms of $\mathcal{A}$ that conjugate by $v_{i}$, then the based partition associated to $v_{i}$ in $\tau(\mathcal{A})$ will be $\tau(\mathcal{A})_{i}=\{\{v_{i}\},I_{1},...,I_{k},M\}$; where $M=V(\Gamma)-(\mathrm{st}(v_{i})\cup I_{1}\cup\cdots\cup I_{k})$, and $M$ contains the minimal element of $\Gamma-\mathrm{st}(v_{i})$. It is easy to check that for all vertex types $\tau$ the cone point of its canonical basis is $\tau$ itself; i.e., $\tau(\mathcal{B}(\tau))=\tau$. And for all compatible collections $\mathcal{A}$, the canonical basis of its cone point is $\mathcal{A}$ itself; i.e., $\mathcal{B}(\tau(\mathcal{A}))=\mathcal{A}$. As a consequence, the cone point $\tau(\mathcal{A})$ lies in $\mathrm{supp}(\mathcal{A})$.
@@
5. Computing $H^{*}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$
We want to compute the cohomology groups $H^{q}\left(\Sigma\mathrm{POut}(A_{\Gamma})\right)$ using the equivariant spectral sequence associated to the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on $\mathrm{MM}_{\Gamma}$. Recall that if $G$ is a group acting simplicially on a contractible complex $X$, the first page of this equivariant spectral sequence (see, for example, Proposition 7.3 in Chapter VII of [4]) is

### 16 Apr 2026 v2

# Cohomology of pure symmetric automorphisms of right-angled Artin groups

@@ Abstract.
We compute the cohomology groups of the pure symmetric outer automorphism group $\Sigma\mathrm{POut}(A_{\Gamma})$ and the pure symmetric automorphism group $\Sigma\mathrm{PAut}(A_{\Gamma})$ of a right-angled Artin group $A_{\Gamma}$. Using the equivariant spectral sequence arising from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on the generalized McCullough-Miller complex $\mathrm{MM}_{\Gamma}$, we show that $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is free abelian and we compute its rank in terms of the combinatorics of certain poset. Applying the Lyndon-Hochschild-Serre spectral sequence and the Leray-Hirsch theorem we do the same for $H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))$. In both cases the cohomology algebra is generated in degree 1. Finally, we introduce the Generalized Brownstein-Lee Conjecture, proposing a presentation of $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$, and prove that it holds in dimension $2$.
@@ 1. Introduction
In [15], McCullough and Miller constructed a contractible simplicial complex $\mathrm{MM}_{n}$ that admits an action of the group $\Sigma\mathrm{POut}(F_{n})$ with free abelian stabilizers. The construction is based on the Whitehead poset $\mathrm{Wh}_{n}$, whose elements can be expressed as labeled trees. In [9], Jensen, McCammond and Meier use the McCullough-Miller complex to compute $H^{*}(\Sigma\mathrm{POut}(F_{n}))$ and $H^{*}(\Sigma\mathrm{PAut}(F_{n}))$, the cohomology groups of the (outer) pure symmetric automorphisms of the free group; even more, they prove the Brownstein-Lee conjecture [5], which stablish a presentation of the cohomology ring $H^{*}(\Sigma\mathrm{PAut}(F_{n}))$.
In [1], Ardaiz-Gale and Martínez-Pérez generalized both the Whitehead poset and the construction of McCullough and Miller from the case of $\Sigma\mathrm{POut}(F_{n})$ to $\Sigma\mathrm{POut}({A_{\Gamma}})$. This new complex, $\mathrm{MM}_{\Gamma}$, admits an action of the group $\Sigma\mathrm{POut}({A_{\Gamma}})$ with free abelian stabilizers and it is also contractible. The powerful combinatorial machinery derived from the labeled trees in $\mathrm{Wh}_{n}$ is the key to prove the main results in [9]. However, vertices in $\mathrm{Wh}_{\Gamma}$ cannot be expressed as labeled trees in general, only as families of based partitions, so most arguments need to be adapted.
In this paper we will generalize some combinatorial results from $\mathrm{Wh}_{n}$ to $\mathrm{Wh}_{\Gamma}$ and use the equivariant spectral sequence derived from the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action over $\mathrm{MM}_{\Gamma}$ to compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ and $H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))$. We show (see Sections 2 and 4 for definitions):
@@
Besides, the cohomology ring $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\bar{\gamma}_{I}^{j}$, where the $C_{I}^{j}$ are canonical automorphisms.
@@
$H^{q}(\Sigma\mathrm{PAut}(A_{\Gamma}))\cong\bigoplus_{i+j=q}\mathbb{Z}^{K_{i}\cdot N_{j}}.$
Besides, the cohomology ring $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated in degree 1.
Furthermore, we generalize the relations from the Brownstein-Lee Conjecture to the case of RAAGs, formulate the Generalized Brownstein-Lee Conjecture, and prove that it holds in dimension 2.
Conjecture A (Generalized Brownstein-Lee Conjecture).
Let $\gamma_{A}^{i}$ be a one-dimensional class in $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$, where $A$ is a connected component in $\Gamma-\mathrm{st}(v_{i})$. For each pair $v_{i},v_{j}$ of two non-adjacent vertices in $\Gamma$, let $D^{j},D^{i}$ be the respective dominant components and $\mathcal{C}$ a shared component. Then $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ is generated by the one-dimensional classes $\gamma_{A}^{i}$, modulo the following relations for each shared component $\mathcal{C}$ of $v_{i},v_{j}$.
@@
$\phi:H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))\twoheadrightarrow R$
@@
The paper is organized as follows: In Section 2 we explain some basic results about MMΓ, obtained from [1]. In Section 3, we go over the main definitions in [9]. In Section 4 we generalize the notion of one-two automorphism which plays a key role in [9] and do some technical work that will be needed later. In Section 5 we apply the previous results to the spectral sequence derived from the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on MMΓ to compute $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$ and prove Theorem A. In Section 6, we use the Lyndon-Hoschild-Serre spectral sequence associated to $1\to A_{\Gamma}\to\Sigma\mathrm{PAut}(A_{\Gamma})\to\Sigma\mathrm{POut}(A_{\Gamma})\to 1$ and the Leray-Hirsch Theorem to compute the $H^{*}(\Sigma\mathrm{PAut}(A_{\Gamma}))$ groups and prove Theorem B. In Section 7 we take the first steps into proving the Brownstein-Lee conjecture for $\Sigma\mathrm{PAut}(A_{\Gamma})$ and we prove Theorem C.
@@ 2. The McCullough-Miller complex
We will use $\Gamma$ to denote a finite simplicial graph with vertex set $V(\Gamma)$ and set of edges $E(\Gamma)$. Very often we will identify $\Gamma$ or a subgraph of $\Gamma$ with its vertex set. Let $v\in V(\Gamma)$, $\mathrm{lk}(v)$ denotes the link of $v$, which is the subgraph induced by the vertices adjacent to $v$. We will define the star of $v$, $\mathrm{st}(v)$, as the subgraph induced by $\mathrm{lk}(v)\cup\{v\}$.
@@ 2.1. Pure symmetric automorphisms of RAAGs
Given $v\in L=V(\Gamma)\cup V(\Gamma)^{-1}$ and $A$ a union of connected components of $\Gamma-\mathrm{st}(v)$, the partial conjugation of $A$ by $v$ is the automorphism $C_{A}^{v}$ given by
@@
The group $\Sigma\mathrm{PAut}(A_{\Gamma})$ is the group generated by all partial conjugations $C_{A}^{u}$ for $u\in\Gamma$ and $A$ a connected component of $\Gamma-\mathrm{st}(u)$, subject to the following relations:
@@
The nuclear vertex type, denoted $\mathcal{O}$, is the vertex type such that for each $v\in\Gamma$ the associated based partition is trivial, i.e., $\mathcal{O}_{v}=\{\{v\},P\}$, where $P=\Gamma-\mathrm{st}(v)$. It is the unique minimal element in the $\mathrm{Wh}_{\Gamma}$ poset.
@@
We say that a generating system $X$ of $A_{\Gamma}$ is a basis if there is some $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$ such that $X=\alpha(V(\Gamma))$. A marked based partition is a pair $(X,\tau_{u})$ where $X$ is an ordered basis of $A_{\Gamma}$ and $\tau_{u}$ is a based partition and a marked vertex type $(X,\tau)$ is the family of marked based partitions $(X,\tau_{u})$ where $\tau_{u}$ runs over the set of based partitions of a vertex type $\tau$. By marking a based partition $\tau_{u}$ we understand that we are relabeling the elements in the partition using $X$ (via the automorphism $\alpha$). We extend the partial order of Definition 2.11 to marked vertex types as follows: $(X_{1},\tau)\leq(X_{2},\tau^{\prime})$ if $X_{1}=X_{2}$ and $\tau\leq\tau^{\prime}$. The marked vertex types form a poset and the group $\Sigma\mathrm{PAut}(A_{\Gamma})$ acts on this poset via
@@
where $\alpha\in\Sigma\mathrm{PAut}(A_{\Gamma})$. We say that $\alpha$ is carried by a marked vertex type $(X,\tau)$ if $\gamma\alpha\gamma^{-1}$ is carried by $\tau$ where $\gamma(X)=V(\Gamma)$.
@@
The definition of the action above implies that the stabilizer $H=\mathrm{Stab}_{\Sigma\mathrm{POut}({A_{\Gamma}})}[V(\Gamma),\tau]$ of the vertex $V(\Gamma),\tau]$ under the $\Sigma\mathrm{POut}({A_{\Gamma}})$ action is precisely the subgroup generated by those automorphisms carried by the vertex type $\tau$. This implies that the stabilizer of the vertex $X,\tau]$ is $\alpha H\alpha^{-1}$ where $\alpha$ is such that $X=\alpha(V(\Gamma))$. It is proved in [1, Theorem A] that these stabilizers are free abelian, that their rank can be computed in terms of the length of the based partitions of $\tau$ and that they yield in fact all the cell stabilizers of the McCullough-Miller complex $\mathrm{MM}_{\Gamma}$.
@@
$\{\mathrm{Stab}_{\Sigma\mathrm{POut}(A_{\Gamma})}[V(\Gamma),\tau]\mid\tau\text{ vertex type}\}$
@@ 4. Choosing the basis for $\Sigma\mathrm{POut}({A_{\Gamma}})$
From now on we fix an order in the elements in $V(\Gamma)$, and denote them by $v_{1},...,v_{n}$. Sometimes, for simplicity and in the aim of a clear notation, we will just identify $v_{i}$ with $i$; in this sense, we will often put $\tau_{i}$ to denote the based partition $\tau_{v_{i}}$, or $C^{j}_{I}$ for the partial conjugation $C^{v_{j}}_{I}$. Let $\tau$ be a vertex type in MMΓ and $\tau_{i}$ a based partition of $\tau$. The product of the partial conjugations associated to each petal of $\tau_{i}$ gives the inner automorphism associated to $v_{i}$, so it is trivial in the outer group. We need to have a way to choose families of partial conjugations associated to $\tau_{i}$ which are independent.
@@
Following the definition, we have that if $\bar{C}_{I_{1}}^{i},...,\bar{C}_{I_{k}}^{i}$ is the set of canonical outer-automorphisms of $\mathcal{A}$ that conjugate by $v_{i}$, then the based partition associated to $v_{i}$ in $\tau(\mathcal{A})$ will be $\tau(\mathcal{A})_{i}=\{\{v_{i}\},I_{1},...,I_{k},M\}$; where $M=\Gamma-(\mathrm{st}(v_{i})\cup I_{1}\cup\cdots\cup I_{k})$, and $M$ contains the minimal element of $\Gamma-\mathrm{st}(v_{i})$. It is easy to check that for all vertex types $\tau$ the cone point of its canonical basis is $\tau$ itself; i.e., $\tau(\mathcal{B}(\tau))=\tau$. And for all compatible collections $\mathcal{A}$, the canonical basis of its cone point is $\mathcal{A}$ itself; i.e., $\mathcal{B}(\tau(\mathcal{A}))=\mathcal{A}$. As a consequence, the cone point $\tau(\mathcal{A})$ lies in $\mathrm{supp}(\mathcal{A})$.
@@
5. Computing $H^{*}(\Sigma\mathrm{POut}(A_{\Gamma}))$
We want to compute the cohomology groups $H^{q}(\Sigma\mathrm{POut}(A_{\Gamma}))$ using the equivariant spectral sequence associated to the action of $\Sigma\mathrm{POut}(A_{\Gamma})$ on $\mathrm{MM}_{\Gamma}$. Recall that if $G$ is a group acting simplicially on a contractible complex $X$, the first page of this equivariant spectral sequence (see, for example, Proposition 7.3 in Chapter VII of [4]) is